« Project Status: 4 January 2008 2030UTC | Main | Project Status: 6 January 2008 2030UTC »

Project Status: 5 Jan 2008 2030UTC

IOP #4 Still Producing Precipitation

Weather Status: Rain and snow was quite heavy over and near the ARB. BLU's power trouble gave some uncertainty in estimating the amount of rain and snow. Reports from road interests indicated 2-3 feet of snow, perhaps 4 feet; liquid was more uncertain but Huysink had a 24 hour total of 5.44 inches ending at 5 Jan 12UTC (4L), and it was still snowing at that time. One report from a nearby mountain ridge had 9 inches of liquid, so snow in favored places could have gone up to 10 feet! As stated yesterday IOP#4 certainly lived up to expectations.....and it's not over. A series of waves imbedded in the cold air of the parent trough will continue to pinwheel into the ARB through early Monday. A wave today that produced another round of heavy rain on the coast is now moving into the Sierras. The NAM and high resolution ensembles are indicating another 3-4 inches of liquid by Monday morning (7 Jan 12UTC [04L]) from this current minor wave and the two others that will follow. One ensemble member even suggested over 5 inches. So it appears that IOP#4 will continue through Monday morning when NWP suggests there will be a clean break. Following the break another system will move into the ARB beginning early Tuesday (8 Jan 12UTC [04L]), starting with a warm advection event and another IPW plume. This event will be short lived ending late Tuesday (9 Jan 06UTC [22L]). Freezing levels will start low (5000ft) rise a bit but then drop, so the HYDROX site may be in the melting layer at best and will not see a pure rain event. Another completely separate system is due to arrive late Wednesday (10 Jan 03UTC [19L]). This system too, will have an associated IPW plume. If these pan out we may have to declare IOP#5 and 6.

Summary: IOP#4 precipitation continuing through early Monday. New system moves in early Tuesday (8 Jan 12UTC [04L]) through late Tuesday (9 Jan 06UTC [8/22L]...outlook is 1.5-2inches). A second trough moves in late Wednesday (10 Jan 06UTC [9/22L])ending early Thursday.

Observational Status: Power or telecommunications problems have taken out surface observations at BLU and Big Bend, and surface and profiler observations at Cazadero and Chico. Slough House's power came back on Friday afternoon (5 Jan 00UTC [16L[). SHS soundings were launched pretty much on schedule at 5/00, 04, 08, and 12UTC. The 12UTC sondes at RNO, OAK and SHS were the last of IOP#4. With the impending troughs on the horizon we can tentatively put crews at SHS, OAK, and RNO for supplemental sondes on alert for early Tuesday. Likewise the HYDROX radar crew will likely redeploy as soon as access to the site is restored on Monday. Operations may have to commence late Monday (8 Jan 06UTC [7/22L]) or early Tuesday. Forecasts will allow these start times to be better defined at tomorrow's telecon.

Decision Summary: IOP#4 continues through early Monday (7 Jan 12UTC [04L]). No special observations required (radar and sondes). Radar and sonde crews put on alert for possible deployment and initiation late Monday or early Tuesday. This will be better specified on Sunday. A decision will also be made on a radar crew change based upon a sufficiently long break between the Tuesday and Wednesday events.

Product Status: Products from HMT affected observing sites are down owing to power or telecommunications problems. High resolution model runs are complete and available through the HMT web page.

McGinley GSD

TrackBack

TrackBack URL for this entry:
http://www.etl.noaa.gov/cgi-bin/mt/mt-tb.cgi/657